From 200mg of co2,10^21 co2 molecules are removed . Calculate the weight of the remaining co2.

Grade:12

1 Answers

Arun
6 years ago
 
No of moles of CO2 initially = weight of substance(gm)/ molecular mass
No.of moles of CO2 initially = 0.2/44 = 1/220 mol
No. of moles removed = no. of molecules/ avagardo constant
No. of moles removed = 1021/ (6 × 1023) = 1/600 mol
No. of moles left = 1/220 – 1/600 = 0.028 (approx.)